# PDLP: A Practical First-Order Method for Large-Scale Linear Programming

### v1

@@ 3 Practical algorithmic improvements
Note that while our enhancements are inspired by theory, our focus is on practical performance. The algorithm as a whole does not have convergence guarantees. Guarantees for individual enhancements are noted at the end of the section.
@@
There are some minor differences between the pseudo-code and the actual code. In particular, we only evaluate the restart or termination criteria (Line 1) every $64$ iterations. This reduces the associated overhead of these evaluations with minimal impact on the total number of iterations. We also check the termination criteria before beginning the algorithm and if we detect a numerical error.
1Input: An initial solution $z^{0,0}$;
2 Initialize outer loop counter $n\leftarrow 0$, total iterations $k\leftarrow 0$, step size $\hat{\eta}^{0,0}\leftarrow 1/\|A\|_{\infty}$, primal weight $\omega^{0}\leftarrow$ InitializePrimalWeight(c, q);
3 repeat
4       $t\leftarrow 0$;
5       repeat
6             $z^{n,t+1},\eta^{n,t+1},\hat{\eta}^{n,t+1}\leftarrow$ AdaptiveStepOfPDHG($z^{n,t},\omega^{n},\hat{\eta}^{n,t},k$) ;
7             $\bar{z}^{n,t+1}\leftarrow\frac{1}{\sum_{i=1}^{t+1}\eta^{n,i}}\sum_{i=1}^{t+1}\eta^{n,i}z^{n,i}$ ;
8             $z_{\text{c}}^{n,t+1}\leftarrow$ GetRestartCandidate($z^{n,t+1},\bar{z}^{n,t+1},z^{n,0}$) ;
9             $t\leftarrow t+1$, $k\leftarrow k+1$ ;
10
11      until restart or termination criteria holds;
12       restart the outer loop. $z^{n+1,0}\leftarrow z_{\text{c}}^{n,t}$, $n\leftarrow n+1$;
13       $\omega^{n}\leftarrow$PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$) ;
14
15until termination criteria holds;
@@ 13       $\omega^{n}\leftarrow$PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$) ;
1Function AdaptiveStepOfPDHG ($z^{n,t}$, $\omega^{n}$, $\hat{\eta}^{n,t},k$):
2       $(x,y)\leftarrow z^{n,t}$, $\eta\leftarrow\hat{\eta}^{n,t}$ ;
3       for $i=1,\dots,\infty$ do
4             $x^{\prime}\leftarrow\mathop{\bf proj}_{\mathcal{X}}(x-\frac{\eta}{\omega^{n}}(c-A^{\top}y))$ ;
5             $y^{\prime}\leftarrow y-A(2x^{\prime}-x)-\eta\omega^{n}\mathop{\bf proj}_{[\ell_{c},u_{c}]}\left((\eta\omega^{n})^{-1}y-A(2x^{\prime}-x)\right)$;
6             $\bar{\eta}\leftarrow\frac{\|(x^{\prime}-x,y^{\prime}-y)\|_{\omega^{n}}^{2}}{2(y^{\prime}-y)^{\top}A(x^{\prime}-x)}$ ;
7             $\eta^{\prime}\leftarrow\min\left((1-(k+1)^{-0.3})\bar{\eta},(1+(k+1)^{-0.6})\eta\right)$ ;
8             if $\eta\leq\bar{\eta}$ then
9                   return $(x^{\prime},y^{\prime})$, $\eta$, $\eta^{\prime}$
10             end if
11            $\eta\leftarrow\eta^{\prime}$ ;
12
13       end for
14
15
@@ 11            $\eta\leftarrow\eta^{\prime}$ ;
The convergence analysis [25, Equation (15)] of standard PDHG (4) relies on a small constant step size
@@
where $z^{k}=(x^{k},y^{k})$. Classically one would ensure (6) by picking $\eta=\frac{1}{\|A\|_{2}}$. This is overly pessimistic and requires estimation of $\|A\|_{2}$. Instead our AdaptiveStepOfPDHG adjusts $\eta$ dynamically to ensure that (6) is satisfied. If (6) isn’t satisfied, we abort the step; i.e., we reduce $\eta$, and try again. If (6) is satisfied we accept the step. This is described in Algorithm 2. Note that in Algorithm 2 $\bar{\eta}\geq\frac{1}{\|A\|_{2}}$ holds always, and from this one can show the resulting step size $\eta\geq\frac{1-o(1)}{\|A\|_{2}}$ holds as $k\rightarrow\infty$.
@@
In PDLP, we adaptively restart the PDHG algorithm in each outer iteration. The key quantity to our restarts at the $n$-th outer iteration is the normalized duality gap at $z$ which for any radius $r\in(0,\infty)$ is defined by
$\rho_{r}^{n}(z):=\frac{1}{r}\max_{(\hat{x},\hat{y})\in\{\hat{z}\in Z:\|\hat{z}-z\|_{\omega^{n}}\leq r\}}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\},$
introduced by [8]. Unlike the standard duality gap, $\max_{(\hat{x},\hat{y})\in Z}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\},$ the normalized duality gap is always a finite quantity. Furthermore, for any value of $r$ and $\omega^{n}$, the normalized duality gap $\rho_{r}^{n}(z)$ is 00 if and only if the solution $z$ is an optimal solution to (3) [8]; thus, it provides a metric for measuring progress towards the optimal solution. The normalized duality gap is computable in linear time [8]. For brevity, define $\mu_{n}(z,z_{\text{ref}})$ as the normalized duality gap at $z$ with radius $\|z-z_{\text{ref}}\|_{\omega^{n}}$, i.e.,
@@
[…]
@@
The motivation for (i) is presented in [8]; it guarantees the linear convergence of restarted PDHG on LP problems. The second condition in (ii) is inspired by adaptive restart schemes for accelerated gradient descent where restarts are triggered if the function value increases [71]. The first inequality in (ii) provides a safeguard for the second one, preventing the algorithm from restarting every inner iteration or never restarting. The motivation for (iii) relates to the primal weights (Section 3). In particular, primal weight updates only occur after a restart, and condition (iii) ensures that the primal weight will be updated infinitely often. This prevents a bad choice of primal weight in earlier iterations from causing progress to stall for a long time.
@@
Define the map Combine$(\ell_{c},u_{c})$ which outputs a vector in $\mathbb{R}^{m}$ whose $i$th coordinate is given by $\max\left\{0,(\ell_{c})_{i}^{\star},(u_{c})_{i}^{\star}\right\}$ where for any $x\in\mathbb{R}\cup\{\pm\infty\}$, the operation $(x)^{\star}$ is equal to $|x|$ if it is finite or equal to zero otherwise. The primal weight is initialized using
@@
where $\epsilon_{\text{zero}}$ is a small nonzero tolerance.
1 Function PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$):
2       $\Delta_{x}^{n}=\|x^{n,0}-x^{n-1,0}\|_{2},\quad\Delta_{y}^{n}=\|y^{n,0}-y^{n-1,0}\|_{2}$ ;
3       if $\Delta_{x}^{n}>\epsilon_{\text{zero}}$ and $\Delta_{y}^{n}>\epsilon_{\text{zero}}$ then
4             return $\exp\left(\theta\log\left(\frac{\Delta_{y}^{n}}{\Delta_{x}^{n}}\right)+(1-\theta)\log(\omega^{n-1})\right)$
5       else
6             return $\omega^{n-1}$ ;
7
8       end if
9
10
Algorithm 3 Primal weight update
Algorithm 3 aims to choose the primal weight $\omega^{n}$ such that distance to optimality in the primal and dual is the same, i.e., $\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}\approx\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}$. By definition of $\|\cdot\|_{\omega}$,
$\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}=\omega^{n}\|x^{n,t}-x^{\star}\|_{2},\quad\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}=\frac{1}{\omega^{n}}\|y^{n,t}-y^{\star}\|_{2}.$
Setting these two terms equal yields $\omega^{n}=\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$. Evidently, the quantity $\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$ is unknown beforehand, but we attempt to estimate it using $\Delta^{n}_{y}/\Delta^{n}_{x}$. However, the quantity $\Delta^{n}_{y}/\Delta^{n}_{x}$ can change wildly from one restart to another, causing $\omega^{n}$ to oscillate. To dampen variations in $\omega^{n}$, we first move to a log-scale where the primal weight is symmetric, i.e., $\log(1/\omega^{n})=-\log(\omega^{n})$, and perform an exponential smoothing with parameter $\theta\in[0,1]$. In PDLP, we use $\theta=0.5$.
There are several important differences between our primal weight heuristic and proposals from Goldstein et al. [40, 41]. For example, Goldstein et al. make relatively small changes to the primal weights at each iteration, attempting to balance the primal and dual residual. These changes have to be diminishingly small because, in our experience, PDHG may be unstable if they are too big. In contrast, in our method, the primal weight is only updated during restarts, which in practice allows for much larger changes without instability issues. Moreover, our scheme tries to balance the weighted distance traveled in the primal and dual rather than the residuals.
@@ 1 Function PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$):
  • Pock-Chambolle [75]: Pock and Chambolle proposed a family of diagonal preconditioners [^3] Diagonal preconditioning is equivalent to changing to a weighted $\ell_{2}$ norm in the proximal step of PDHG (weight defined by $D_{2}$ and $D_{1}$ for the primal and dual respectively). Pock and Chambolle use this weighted norm perspective. for PDHG parameterized by $\alpha$, where the diagonal matrices are defined by $(D_{1})_{jj}=\sqrt{\|A_{j,\cdot}\|_{2-\alpha}}$ for $j=1,...,m_{1}+m_{2}$ and $(D_{2})_{ii}=\sqrt{\|A_{\cdot,i}\|_{\alpha}}$ for $i=1,...,n$. We use $\alpha=1$ in PDLP. This is the baseline diagonal preconditioner in the PDHG literature.
  • Ruiz [80]: In an iteration of Ruiz scaling, the diagonal matrices are defined as $(D_{1})_{jj}=\sqrt{\|A_{j,\cdot}\|_{\infty}}$ for $j=1,...,m_{1}+m_{2}$ and $(D_{2})_{ii}=\sqrt{\|A_{\cdot,i}\|_{\infty}}$ for $i=1,...,n$. Ruiz [80] shows that if this rescaling is applied iteratively, the infinity norm of each row and each column converge to $1$.
For the default PDLP settings, we apply a combination of Ruiz rescaling [80] and the preconditioning technique proposed by Pock and Chambolle [75]. In particular, we apply $10$ iterations of Ruiz scaling and then apply the Pock-Chambolle scaling.
@@
Inspired by [9], PDLP periodically checks whether any of three sequences—the difference of consecutive iterates, the normalized iterates, or the normalized running average—can serve as approximate certificates of infeasibility. We evaluate running averages starting from the most recent restart. Computing these sequences imposes little additional overhead, allowing PDLP to detect infeasibility efficiently. Notably, verifying all three sequences is often faster than relying on just one, since their convergence speed varies with the problem’s geometry [9].
@@
While PDLP’s enhancements are motivated by theory, some of them may not preserve theoretical guarantees: we do not have a proof of convergence for the adaptive step size rule or primal weight update scheme. However, Applegate et al. [8] showed in a simpler setting with fixed primal weight and step size that adaptive restarts preserve convergence guarantees. Similarly, the theoretical gurantees for infeasibility detection [9] do not account for restarts.

### v2

@@ 3 Practical algorithmic improvements
Note that while our enhancements are inspired by theory, our focus is on practical performance. The algorithm as a whole does not have convergence guarantees. Guarantees for certain individual enhancements are noted at the end of the section.
@@
There are some minor differences between the pseudo-code and the actual code. In particular, we only evaluate the restart or termination criteria (Line 1) every $64$ iterations [^3] Condition (iii) of the restart criteria is checked every iteration.. This reduces the associated overhead of these evaluations with minimal impact on the total number of iterations. We also check the termination criteria before beginning the algorithm and if we detect a numerical error.
Input: An initial solution $z^{0,0}$;
Initialize outer loop counter $n\leftarrow 0$, total iterations $k\leftarrow 0$, step size $\hat{\eta}^{0,0}\leftarrow 1/\|A\|_{\infty}$, primal weight $\omega^{0}\leftarrow$ InitializePrimalWeight($c,\ell_{c},u_{c}$);
repeat
 $t\leftarrow 0$;
 repeat
    $z^{n,t+1},\eta^{n,t+1},\hat{\eta}^{n,t+1}\leftarrow$ AdaptiveStepOfPDHG($z^{n,t},\omega^{n},\hat{\eta}^{n,t},k$) ;
    $\bar{z}^{n,t+1}\leftarrow\frac{1}{\sum_{i=1}^{t+1}\eta^{n,i}}\sum_{i=1}^{t+1}\eta^{n,i}z^{n,i}$ ;
    $z_{\text{c}}^{n,t+1}\leftarrow$ GetRestartCandidate($z^{n,t+1},\bar{z}^{n,t+1},z^{n,0}$) ;
    $t\leftarrow t+1$, $k\leftarrow k+1$ ;
 until restart or termination criteria holds;
 restart the outer loop. $z^{n+1,0}\leftarrow z_{\text{c}}^{n,t}$, $\hat{\eta}^{n+1,0}\leftarrow\hat{\eta}^{n,t}$, $n\leftarrow n+1$ ;
 $\omega^{n}\leftarrowPrimalWeightUpdate$($z^{n,0},z^{n-1,0},\omega^{n-1}$) ;
until termination criteria holds;
@@
Function AdaptiveStepOfPDHG ($z^{n,t}$, $\omega^{n}$, $\hat{\eta}^{n,t},k$):
 $(x,y)\leftarrow z^{n,t}$, $\eta\leftarrow\hat{\eta}^{n,t}$ ;
 for $i=1,\dots,\infty$ do
    $x^{\prime}\leftarrow\mathop{\bf proj}_{\mathcal{X}}(x-\frac{\eta}{\omega^{n}}(c-A^{\top}y))$ ;
    $y^{\prime}\leftarrow y-\eta\omega^{n}A(2x^{\prime}-x)-\eta\omega^{n}\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}\left((\eta\omega^{n})^{-1}y-A(2x^{\prime}-x)\right)$;
    $\bar{\eta}\leftarrow\begin{cases}\frac{\|(x^{\prime}-x,y^{\prime}-y)\|_{\omega^{n}}^{2}}{2(y^{\prime}-y)^{\top}A(x^{\prime}-x)}&2(y^{\prime}-y)^{\top}A(x^{\prime}-x)>0\\ \infty&\text{otherwise}\end{cases}$ ;
    $\eta^{\prime}\leftarrow\min\left((1-(k+1)^{-0.3})\bar{\eta},(1+(k+1)^{-0.6})\eta\right)$ ;
    if $\eta\leq\bar{\eta}$ then
       return $(x^{\prime},y^{\prime})$, $\eta$, $\eta^{\prime}$
      end if
    $\eta\leftarrow\eta^{\prime}$ ;
   end for
@@
The convergence analysis [27, Equation (15)] of standard PDHG (3) relies on a small constant step size
@@
where $z^{k}=(x^{k},y^{k})$. Classically one would ensure (5) by picking $\eta=\frac{1}{\|A\|_{2}}$. This is overly pessimistic and requires estimation of $\|A\|_{2}$. Instead our AdaptiveStepOfPDHG adjusts $\eta$ dynamically to ensure that (5) is satisfied. If (5) isn’t satisfied, we abort the step; i.e., we reduce $\eta$, and try again. If (5) is satisfied we accept the step. This is described in Algorithm 2. Note that in Algorithm 2 $\bar{\eta}\geq\frac{1}{\|A\|_{2}}$ holds always, and from this one can show the step size $\eta$ satisfies $\eta\geq\frac{1-o(1)}{\|A\|_{2}}$ as $k$ tends to infinity.
This step size is compared against other standard step size schemes in Appendix C.1 of Applegate et al. [9].
@@
In PDLP, we adaptively restart the PDHG algorithm at each outer iteration. To decide when to restart at the $n$th outer iteration we use the normalized duality gap at $z$, which for any radius $r\in(0,\infty)$, is defined by
$\rho_{r}^{n}(z):=\frac{1}{r}\max_{(\hat{x},\hat{y})\in\{\hat{z}\in Z:\|\hat{z}-z\|_{\omega^{n}}\leq r\}}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\}$
where $Z=\mathcal{X}\times\mathcal{Y}$ and for completeness define $\rho_{0}(z):=\limsup_{r\rightarrow 0^{+}}\rho_{r}^{n}(z)$. This measure was introduced by Applegate et al. [10]. Unlike the standard duality gap, $\max_{(\hat{x},\hat{y})\in Z}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\},$ the normalized duality gap is always a finite quantity. Furthermore, for any value of $r$ and $\omega^{n}$, the normalized duality gap $\rho_{r}^{n}(z)$ is 0 if and only if the solution $z$ is an optimal solution to (2) [10]; thus, it provides a metric for measuring progress towards the optimal solution. We compute the normalized duality gap based on the linear time implementation given by Applegate et al. [10, Section 3.2]. For brevity, define $\mu_{n}(z,z_{\text{ref}})$ as the normalized duality gap at $z$ with radius $\|z-z_{\text{ref}}\|_{\omega^{n}}$, i.e.,
@@
$\text{{{GetRestartCandidate(}}{\emph{$z^{n,t+1},\bar{z}^{n,t+1},z^{n,0}$}}{{)}}}:=\begin{cases}z^{n,t+1}&\mu_{n}(z^{n,t+1},z^{n,0})<\mu_{n}(\bar{z}^{n,t+1},z^{n,0})\\ \bar{z}^{n,t+1}&\text{otherwise}\ .\end{cases}$
@@
The motivation for (i) is presented in [10]; it guarantees the linear convergence of restarted PDHG on LP problems. The second condition in (ii) is inspired by adaptive restart schemes for accelerated gradient descent where restarts are triggered if the function value increases [78]. The first inequality in (ii) provides a safeguard for the second one, preventing the algorithm from restarting every inner iteration or never restarting. The motivation for (iii) relates to the primal weight updates. In particular, primal weight updates only occur after a restart (see Line 1 of Algorithm 1), and condition (iii) ensures that the primal weight will be updated infinitely often. This prevents a bad choice of primal weight in earlier iterations from causing progress to stall for a long time.
The practical improvement from our restart scheme is demonstrated in Appendix C.2 of Applegate et al. [9].
@@
Algorithm 3 aims to choose the primal weight $\omega^{n}$ such that distance to optimality from primal and dual iterates is the same, i.e., $\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}\approx\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}$. By definition of $\|\cdot\|_{\omega}$,
$\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}=\sqrt{\omega^{n}}\|x^{n,t}-x^{\star}\|_{2},\quad\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}=\frac{1}{\sqrt{\omega^{n}}}\|y^{n,t}-y^{\star}\|_{2}.$
Setting these two terms equal yields $\omega^{n}=\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$. Evidently, the quantity $\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$ is unknown beforehand, but we attempt to estimate it using $\Delta^{n}_{y}/\Delta^{n}_{x}$. However, the quantity $\Delta^{n}_{y}/\Delta^{n}_{x}$ can change wildly from one restart to another, causing $\omega^{n}$ to oscillate. To dampen variations in $\omega^{n}$, we first move to a log-scale, in which the primal weight is symmetric, i.e., $\log(1/\omega^{n})=-\log(\omega^{n})$, and perform an exponential smoothing with parameter $\theta\in[0,1]$. In PDLP, we use $\theta=0.5$. The parameter $\epsilon_{\text{zero}}=10^{-10}$ is a small nonzero tolerance that, in principle, could be adjusted. It helps prevent numerical issues or other performance issues occurring from very huge or very small movement of the iterates resulting in a very dramatic change in the primal weights.
Function PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$):
 $\Delta_{x}^{n}=\|x^{n,0}-x^{n-1,0}\|_{2},\quad\Delta_{y}^{n}=\|y^{n,0}-y^{n-1,0}\|_{2}$ ;
 if $\Delta_{x}^{n}\in(\epsilon_{\text{zero}},\epsilon_{\text{zero}}^{-1})$ and $\Delta_{y}^{n}\in(\epsilon_{\text{zero}},\epsilon_{\text{zero}}^{-1})$ then
    return $\exp\left(\theta\log\left(\frac{\Delta_{y}^{n}}{\Delta_{x}^{n}}\right)+(1-\theta)\log(\omega^{n-1})\right)$
 else
    return $\omega^{n-1}$ ;
   end if
Algorithm 3 Primal weight update
Initialization of the primal weights.
@@
$\textnormal{{InitializePrimalWeight}}(c,\ell_{c},u_{c}):=\begin{cases}\frac{\|c\|_{2}}{\|\texttt{Combine}(\ell_{c},u_{c})\|_{2}}&\|c\|_{2}>0.0\text{ and }\|\texttt{Combine}(\ell_{c},u_{c})\|_{2}>0.0\\ 1&\text{otherwise}.\end{cases}$
There are several important differences between our primal weight heuristic and proposals from Goldstein et al. [45, 46]. For example, Goldstein et al. make relatively small changes to the primal weights at each iteration, attempting to balance the primal and dual residual. These changes have to be diminishingly small because, in our experience, PDHG may be unstable if they are too big. In contrast, in our method, the primal weight is only updated during restarts, which in practice allows for much larger changes without instability issues. Moreover, our scheme tries to balance the weighted distance traveled by the primal and dual iterates rather than the residuals.
The practical improvement from our primal weight scheme is demonstrated in Appendix C.3 of Applegate et al. [9].
@@
  • Pock-Chambolle [82]: Pock and Chambolle proposed a family of diagonal preconditioners [^4] Diagonal preconditioning is equivalent to changing to a weighted $\ell_{2}$ norm in the proximal step of PDHG (weight defined by $D_{2}$ and $D_{1}$ for the primal and dual respectively). Pock and Chambolle use this weighted norm perspective. for PDHG parameterized by $\alpha$, where the diagonal matrices are defined by $(D_{1})_{jj}=1/\sqrt{\|A_{j,\cdot}\|_{2-\alpha}}$ for j=1,,mj=1,...,m and $(D_{2})_{ii}=1/\sqrt{\|A_{\cdot,i}\|_{\alpha}}$ for i=1,,ni=1,...,n. We use $\alpha=1$ in PDLP. This is the baseline diagonal preconditioner in the PDHG literature.
  • Ruiz [87]: In an iteration of Ruiz scaling, the diagonal matrices are defined as $(D_{1})_{jj}=1/\sqrt{\|A_{j,\cdot}\|_{\infty}}$ for j=1,,mj=1,...,m and $(D_{2})_{ii}=1/\sqrt{\|A_{\cdot,i}\|_{\infty}}$ for i=1,,ni=1,...,n. Ruiz [87] shows that if this rescaling is applied iteratively, the infinity norm of each row and each column converge to 11.
For the default PDLP settings, we apply a combination of Ruiz rescaling [87] and the preconditioning technique proposed by Pock and Chambolle [82]. In particular, we apply $10$ iterations of Ruiz scaling and then apply the Pock-Chambolle scaling. Appendix C.5 of Applegate et al. [9] shows that PDLP’s diagonal preconditioning outperforms either Pock-Chambolle or Ruiz individually.
@@
Inspired by [11], PDLP periodically checks whether any of three sequences—the difference of consecutive iterates, the normalized iterates, or the normalized running average—can serve as approximate certificates of infeasibility. We evaluate running averages starting from the most recent restart. This allows PDLP to detect infeasibility with little additional overhead. Notably, verifying all three sequences is often faster than relying on just one, since their convergence speed varies with the problem’s geometry [11].
@@
While PDLP’s enhancements are motivated by theory, some of them may not preserve theoretical guarantees: we do not have a proof of convergence for the adaptive step size rule or primal weight update scheme. However, Applegate et al. [10] showed in a simpler setting with fixed primal weight and step size that adaptive restarts preserve convergence guarantees. Similarly, the theoretical guarantees for infeasibility detection [11] do not account for restarts.